how do i solve this write an equation of the line that passes through the given point and is parallel to the given line. (5,4),y=3x+5 and the other one is (5,11),m=3
Answer by drk(1908)

Step 1 - look at the equation and make sure it is expressed as y = mx + b. We have y = 3X + 5.
step 2 - find the slope and create parallel slope. In this equation the slope is 3. A slope parallel to 3 is still 3.
step 3 - take the slope from step 2 and the point (5,4) and put both into y = mx+ b to find b.
4 = 5(3) + b - -> b = -11.
step 4 write the answer. Y = 3X - 11.
---
same steps for second problem:
step 1 - Y = 3x + 5
step 2 - slope = 3; parallel slope = 3.
step 3 - 11 = 5(3) + b - - > b = -4
step 4 - Y = 3X - 4.